How many turns should a solenoid of cross-sectiona
Answer #1

GIVEN:A=3.5*10^-2
l=0.25m
L=40mH

Self inductance:

L=\mu0N^2 A/l

N=sqrt((L*l)/\mu0A)=sqrt((40*10^-3*0.25)/(4*pi*10^-7*3.5*10^-2))=477